the normal boiling point of ethanol is 78.3 deg C and its molar enthalpy of vaporization is 38.56 Kj/mol. what is the change in entropy in the system in J/k when 97.2 grams of ethanol at 1 atm condenses to a liquid at the normal boiling point?
Answer #1

Boiling point Tb = 78.3oC = 351.45 K

ΔH = 38.56 kJ/mol = 38560 J/mol

At the boiling point, ΔG = ΔH - TbΔS = 0

ΔS = ΔH/Tb = 38560/351.45 = 109.7 J/mol.K

Moles of ethanol = mass/molar mass

= 97.2/46.07 = 2.110 mol

Entropy change = moles x ΔS

= 2.110 x 109.7 = 231 J/K